OpenStudy (anonymous):

Help with dividing algebraic fractions? x^3/x-1 divided by x^2/y-xy

OpenStudy (anonymous):

a/b divided by c/d is a/b times d/c

OpenStudy (anonymous):

after writing in this form (multiplication) cancel factors

OpenStudy (jiteshmeghwal9):

\[\dfrac{x^3}{(x-1)} \div \dfrac{x^2}{(y-xy)}\]\[\dfrac{x^3}{(x-1)} \times \dfrac{y(x-1)}{x^2}\]

OpenStudy (jiteshmeghwal9):

This will help u a lot & i think u can do the remaining work now :)

OpenStudy (anonymous):

\[\huge\frac{ x^3 }{ x-1 }\times \frac{ y(x-1) }{ x^2 }\]

OpenStudy (anonymous):

Thank you! (: and now I cancel out the factors...?

OpenStudy (jiteshmeghwal9):

yup & yw :)

OpenStudy (anonymous):

I got xy? Is this correct?